How many different ways can a group of 3 dogs and 2 cats be chosen if a pet store
has 10 dogs and 8 cats?

Answer by vasumathi(46)

Solution :
there are 10 dogs and 8 cats in the store
3 dogs can be chosen in C(10,3)= 120 ways
and 2 cats can be chosen in C(8,2)= 28 ways
then ..
the number of ways in which 3 dogs and 2 cats be chosen is as follows...
C(10,3)*C(8,2)= 120 * 28 =3360